You are considering a 15-year, $1,000 par value bond. Its coupon rate is 8%, and interest is paid semiannually. If you require an "effective" annual interest rate (not a nominal rate) of 8.05%, how much should you be willing to pay for the bond? Do not round intermediate steps. Round your answer to the nearest cent.
Answer #1
face value = 1000
Years remaining to Maturity = 15
Semiannual periods (n)= (15*2) = 30
Coupon rate = 8.00%
Semiannual Coupon = 1000*8%/2 = 40
Required annual effective rate = 8.05%

EAR = ((1+semiannual rate)^no of compounding in year)-1

8.05% =( (1+i)^2)-1

1+0.0805 = (1+i)^2

1+i= (1.0805)^(1/2)= 1.039471019
i= 0.03947101932
so Semiannual yield = 0.03947101932

Bond price formula = Coupon amount * (1 - (1/(1+i)^n)/i + face value/(1+i)^n

40*(1-(1/(1+0.03947101932)^30))/0.03947101932 + 1000/(1+0.039471019)^30

1009.206189

So we will be willing to pay for bond $1,009.21

Please thumbs up